# WiLLM: An Open Wireless LLM Communication System

### v1

@@ 7. Conclusion
WiLLM is the first open-source wireless system for LLM services. Our “Tree-Branch-Fruit” slicing architecture, LLMs deployment in CNs, and cross-layer APIs address limitations in existing wireless infrastructure for LLM delivery. The implementation extends OpenAirInterface with components enabling dynamic resource allocation, universal device compatibility, and multi-slice scheduling while preserving backward compatibility. A dataset of 1.6+ million synchronized metrics provides empirical foundation for research, while our smart glasses case demonstrates viability in resource-constrained environments. WiLLM connects telecom evolution with AI, enabling wireless LLM service chain optimization. Future research efforts will focus on addressing the current system’s deficiencies by prioritizing improvements in system stability and fault recovery, ensuring strict compliance with protocols, and implementing security measures for cross-layer APIs.
@@ Appendix A Appendix Overview
This appendix provides supplementary information, analysis, and technical details to support the main body of the paper. The contents are organized as follows: Appendix B presents a general comparison of WiLLM against existing open-source wireless communication systems and LLM deployment frameworks. This comparative analysis establishes WiLLM’s distinctive contributions within the telecommunications-AI convergence landscape.
@@
Appendix F contains supplementary data analysis and visualizations from the WiLLM dataset. These figures provide deeper insights into the stability of experimental conditions, comparative performance between slice-enabled and normal traffic, and the complex interactions between various system parameters. More importantly, we conducted a causal analysis of the results shown in Figure. 6 to Figure  10 of the dataset.
@@
Appendix B Appendix for Section 2 (Related Work and Comparison): General Comparation of Systems
@@
Appendix C Appendix for Section 2 (Related Work and Comparison): Detailed Technical Comparison of Systems
@@ Appendix F Appendix for Section 5 (Dataset)
F.1. Analysis of Uplink vs. Downlink Transmission Patterns (Figures 6, 7 and 8)
The differential performance profiles across slice configurations illustrated in Figure 8 substantiate the efficacy of WiLLM’s “Tree-Branch-Fruit” slicing architecture, with differences primarily attributable to input modality variations. The pronounced disparity in inference time contribution (11.6% to 59.6%) stems fundamentally from the computational asymmetry between processing graphical inputs in uplink scenarios versus textual inputs in downlink scenarios. Image-based queries in uplink transmissions necessitate substantially higher computational overhead for visual feature extraction, semantic interpretation, and multimodal reasoning, consequently dominating the latency profile. Conversely, text-based inputs in downlink scenarios require comparatively less computational intensity, shifting the performance bottleneck toward transmission components (28.9% to 54.7% variation in transmission time). This input-dependent computational variance underscores the necessity of WiLLM’s specialized slicing approach, which adaptively allocates resources based on modality-specific requirements rather than uniform distribution models. The statistically significant performance differentials between slices confirm that WiLLM effectively maintains service isolation while accommodating the inherent computational asymmetry between visual and textual processing pipelines, validating the fundamental premise of modality-optimized “Fruit Slices” within the hierarchical resource management framework.
It is important to note that our current implementation does not utilize specialized image generation models, which would likely introduce additional computational demands in the downlink scenario. Such models might shift the bottleneck toward inference processing rather than transmission, potentially altering the observed latency distribution patterns. Nevertheless, the empirical evidence clearly demonstrates that downlink image transmission presents a more significant bandwidth challenge than uplink image transmission, likely due to differences in compression efficiency and quality requirements between input and output images.
F.2. Resource Allocation-Utilization Relationship Analysis (Figures 9 and 10)
The non-linear correlation between PRB allocation patterns (Figure 9) and actual byte transmission volumes (Figure 10) reveals critical insights into resource utilization efficiency within wireless LLM service delivery. While WiLLM successfully implements differentiated PRB allocation conforming to slice policies, the observed discrepancy between allocated resources and realized throughput reflects complex cross-layer interdependencies spanning the entire system architecture. This phenomenon stems from multi-dimensional factors: (1) variable UE request patterns and buffer state fluctuations that create unpredictable traffic bursts; (2) server-side processing variability due to LLM inference time inconsistencies and computational resource contention; and (3) system execution mechanisms within the gNB layer, including scheduler implementation details, slice prioritization algorithms, and HARQ management policies. The visualization of slice-distinguished traffic with periodic transmission bursts demonstrates how these factors propagate across architectural boundaries, creating emergent system behaviors that cannot be predicted through isolated layer analysis. This empirical evidence substantiates the necessity of WiLLM’s cross-layer API framework, which enables coordinated adaptation across UE configuration, gNB resource allocation, and server processing prioritization based on observed end-to-end performance characteristics rather than theoretical layer-specific models—establishing a comprehensive feedback system essential for optimizing LLM service delivery in wireless environments.
F.3. Dataset Collection Parameters
F.3.1. UE Configuration Parameters
@@
F.3.2. Network Slicing Configuration

### v2

@@ 7. Conclusion
We presented WiLLM, the first open-source platform for wireless LLM service delivery. We identified three characteristics of LLM services over wireless network: bidirectional asymmetry, computation-communication coupling, and token unpredictability. WiLLM addresses these through the new slicing technology, universal compatibility, and cross-layer APIs, enabling AI researchers to experiment with real cellular networks without wireless expertise. The smart glasses case study validated our approach with stable 2-second multimodal responses. By democratizing wireless LLM research, WiLLM transforms a specialized domain into an accessible field for advancing mobile AI services.
@@ Appendix A Appendix Overview
This appendix provides supplementary information, analysis, and technical details to support the main body of the paper. The contents are organized as follows:
Appendix B presents a general comparison of WiLLM against existing open-source wireless communication systems and LLM deployment frameworks. This comparative analysis establishes WiLLM’s distinctive contributions within the telecommunications-AI convergence landscape.
@@
Appendix F contains supplementary data analysis and visualizations from the WiLLM dataset. These figures provide deeper insights into the stability of experimental conditions, comparative performance between slice-enabled and normal traffic, and the complex interactions between various system parameters. More importantly, we conducted a causal analysis of the results shown in Figure. 3 to Figure  7 of the dataset.
@@
Appendix B Appendix for Section 6 (Related Works): General Comparation of Systems
@@
Appendix C Appendix for Section 6 (Related Works): Detailed Technical Comparison of Systems
@@ Appendix F Appendix for Section 5 (Dataset)
F.1. Dataset Collection Parameters
F.1.1. UE Configuration Parameters
@@
F.1.2. Network Slicing Configuration